JET-COOLED LASER SPECTROSCOPY OF A JAHN-TELLER AND PSEUDO JAHN-TELLER ACTIVE MOLECULE: THE NITRATE RADICAL

\underline{MING-WEI CHEN}, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210; KANA TAKEMATSU, MITCHIO OKUMURA, Arthor Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, California 91125; and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

Well-known as an important intermediate in atmospheric chemistry, the nitrate radical (NO_{3}) has been extensively studied both experimentally and theoretically. The three energetically lowest electronic states (X̃ ^{2}A′_{2}, Ã ^{2}E′′, and B̃ 2E′) are strongly coupled by vibronic interactions and hence it is a textbook molecule for understanding the coupling between nearby potential energy surfaces. Such coupling has been treated in considerable detail theoretically.^{a} However, corresponding experimental characterization of the interaction is much less detailed. The experimental results primarily consist of IR measurements of vibrational transitions in the ground state.^{b}^{c} In addition, the electronically forbidden Ã-X̃ transition has been observed in ambient temperature CRDS studies.^{d} To understand both the Jahn-Teller and pseudo Jahn-Teller coupling in the molecule, further measurements are required with different selection rules and/or higher resolution to resolve the rotational structures of different transitions. In our group, a high-resolution (source Δν ≈ 100 MHz in NIR region), jet-cooled CRDS system^{e} can be applied to rotationally resolve the electronically forbidden Ã-X̃ transition. Furthermore, our high-resolution LIF/SEP system (source Δν ≈ 100 MHz) can provide the direct, rotationally resolved measurements of the B̃-X̃ and B̃-Ã transitions by operating in the LIF and SEP modes respectively. Such data can provide unambiguous spectral assignments in the X̃, Ã and B̃ states.

aJ. F. Stanton, J. Chem. Phys., \underline{126}, 134309 (2007)

bK. Kawaguchi, E. Hirota, T. Ishiwata, and I. Tanaka, J. Chem. Phys., \underline{93}, 951 (1990)

cK. Kawaguchi, T. Ishiwata, E. Hirota, and I. Tanaka, Chem. Phys., \underline{231}, 193 (1998)

dA. Deev, J. Sommar, and M. Okumura, J. Chem. Phys., \underline{122}, 224305 (2005)

eS. Wu, P. Dupré, and T. A. Miller, Phys. Chem. Chem. Phys., \underline{8}, 1682, (2006)